# Combined Fit of Spectrum and Composition for FR0 Radio Galaxy Emitted Ultra-High-Energy Cosmic Rays with Resulting Secondary Photons and Neutrinos

### v1

@@ 6.2 Evolving Fraction Fits
The evolving fraction combined fit results for all five simulated intergalactic media and two EAS models
@@
(EPOS-LHC and Sibyll2.3c) compared to Auger data [27, 28] are shown in Figure 10. The QGSJetII-04 model was excluded from the evolving fraction analysis. The mean log mass number shown in Figure 10a illustrates that the inclusion of 44 nuclei fractions leads to a nearly perfect fit across all configurations due to the large number of parameters.
@@
However, similar to the constant fraction fits 1 nG random fields results having a better fit to the lower energy bins, these fields are the best fit to the two highest energy heavier bins with the evolving fractions. As with the constant fraction fits only the Rand.B configurations sufficiently account for the highest energy bin of the energy spectra as shown in Figure 10b. Continuing the trend observed in constant fraction fits, the superior performance of the longer correlation length 1 nG (Rand.B) field at the highest energies remains surprising, challenging the assumption that FR0s are minor contributors at extreme energies. And again, the CLUES model essentially does not contribute to the highest energy bin.
@@
The best-fit parameters are presented in Table 5 and Figure 11, along with corresponding 1 equivalent confidence intervals around the best fit. Table 6 shows the most probable value of the bootstrap sample distributions using an four-dimensional multivariate kernel density estimation (KDE) and the 1 standard deviations of the KDEs. A 48-dimensional KDE estimation has not been done due to time and computing constraints and the 4D KDE should be a reasonable approximation.
@@
For each EAS model and magnetic field, Table 5 and Table 6 show the total , emission spectral index (), exponential rigidity-dependent cutoff (log10()), particle trajectory cutoff (), and relative energy spectrum normalization (). Figure 11 shows the fitted FR0 emission elemental fractions as a function of energy for all 10 configurations.
@@
As indicated in Table 5, similar to findings from constant fraction fits, the CLUES, Rand.A, and Rand.B magnetic fields consistently deliver the best fit results. The Rand.B field, in particular, does stand out as the optimal fit for both the EPOS-LHC and Sibyll2.3c models, with the Rand.B-SIBYLL configuration achieving the overall best fit (Rand.B-EPOS is a very close second). EPOS-LHC is the best fit for all magnetic field models except Rand.B-EPOS but the difference from SIBYLL is very small. The second-best magnetic field after Rand.B is CLUES.
6.2.1 Energy Spectrum Parameters
@@
Overall, the evolving fraction fit energy spectrum parameters of Table 5 and Table 6 again indicate that both magnetic field strength B, and RMS BRMS, have an impact on the expected cosmic-ray source emission energy spectrum. Generally, the emission spectral index, E-γ, increases with the magnetic field strength B. However, the exception is the Rand.B field model where this trend does not apply. The exponential rigidity-dependent cutoff log10() generally rises with an increase in BRMS, except in the case of the Rand.A field model. Furthermore, for smaller magnetic fields, log10() decreases with a heavier composition, while for larger fields, this trend is reversed, demonstrating the complex interaction between magnetic field strength and particle composition. The trajectory cutoff, , is fairly stable across most fields, with the notable exception of the Dolag field which has a significantly higher BRMS of 11 nG.
@@
Despite the no-field and Dolag configurations consistently showing poorer fits, similar to the constant fraction scenario, the inclusion of a large number of parameters in the evolving fraction fits leads to excellent overall fit quality for all configurations. Therefore, the bounds found for the best fit of these parameters from Table 1, using all configurations, are:
1.
@@
Note, that these bounds are larger than the constant fraction fit bounds due the inclusion of the no-field and Dolag configurations.
@@
For the bootstrap KDE fit most probable parameters from Table 6 the ranges found are:
1.
@@
Compared to the recent Auger result of [37] these parameter values are similar to the reported second minima of their models and the low z source distance model (z 0.02).
@@ 6.2.2 UHECR Composition Parameters
The FR0 emission fractions for the evolving energy dependent (observed) fraction fit of proton, helium, nitrogen, silicon, and iron emitted nuclei are shown in Figure 11. Similar to the constant fraction fits, the emission fractions of protons and helium exhibit instability, with dominance of each varying due to propagation effects, as shown in Figure 5. The combined fraction of light components () demonstrates greater stability across different configurations as shown in Figure 11f. Interestingly, for all configurations the fraction of emitted protons increases with energy above 1019 eV as shown in Figure 11a.
@@
The ranges found for the energy bin average FR0 emission best fit elemental fractions (in %) from Figure 11 are:
1.
@@
6.2.3 Multi-messenger Photons and Neutrinos
The evolving fraction fit cosmogenic integral photon and all-flavor neutrino spectra are shown in Figures 12a and 12b. These results are for the best-fit configuration using the Rand.B 1 nG random magnetic field and Sibyll2.3c-based mass composition, compared to the no-field scenario. Although the Sibyll2.3c hadronic interaction model reconstructs a heavier mass for data relative to other models, the evolving fraction fit still predicts that approximately 66% of the FR0 emission consists of light nuclei.
At lower energies (up to eV), the integral photon flux is compatible with pure proton cosmogenic predictions, shifting towards a theoretical mixed composition at higher energies. Interestingly, the neutrino flux resembles predictions for a pure iron prediction, which may be partially influenced by the simulation constraint of relatively close FR0 sources (z 0.2). Given the wide range of theoretical predictions and current experimental upper limits the FR0 simulated flux is reasonable. Overall, for both photon and neutrino spectra, it can be seen that a magnetic field results in a larger flux, particularly at higher energies, as shown in Figure 13.
@@ 7 Summary & Conclusions
This work presents an in-depth study of FR0 radio galaxies, a prevalent but less luminous class of jetted active galaxies, and their potential emission features as ultra-high-energy cosmic ray (UHECR) sources. Although individually less luminous than more powerful radio galaxies, the abundance of FR0 galaxies may render them significant contributors to the UHECR energy density observed on Earth.
@@
To investigate this hypothesis, extensive CRPropa3 [26] simulations were conducted, modeling isotropic cosmic ray emissions extrapolated from measured FR0 properties (density and luminosity distributions and luminosity-redshift source evolution). These simulations were then fit to the mass composition (expressed in mean log mass number ) and energy spectra of UHECR data from the Pierre Auger Observatory [27, 28].
@@
The results from these combined fits, utilizing two fit models, three EAS models, and five intergalactic magnetic field configurations (both random and structured), indicate that emissions from FR0 galaxies can effectively account for the observed UHECR energy spectrum and composition. The overall compatibility, in addition to the results of our previous study of the capability of FR0 radio galaxies to accelerate cosmic rays up to ultra-high energies [6], supports the viability of FR0s as significant contributors to the observed UHECR flux. Furthermore, the secondary photon and neutrino fluxes predicted from UHECR interactions are well within current observational limits, emphasizing the potential of future multi-messenger observations to further refine our understanding of FR0 galaxies as UHECR sources.
@@
For the most extreme energy bin, the configurations tested are generally below the expected flux (only Dolag and longer correlation length 1 nG Rand.B magnetic field models have overlapping error bars, as shown in Figure 6b). Additionally, over all configurations, there is a small underfitting of the last energy bin of (only Dolag and the smaller correlation length 1 nG Rand.A magnetic field models are outside the error bars as shown in Figure 6a). These two results may allow the possibility of a different extreme-energy heavy nuclei source type to become the dominant UHECR flux contributor.
@@
These findings underscore the importance of considering the effects of intergalactic magnetic fields on UHECR propagation. Results across different magnetic field models indicate their significant influence on the arrival energy spectrum and composition of cosmic rays from FR0 galaxies, as well as on resulting secondary photon and neutrino spectra. Larger magnetic fields result in a greater emission of nuclei heavier than helium - additionally, generally, slightly harder emission spectra ( 2.4 compatible with Fermi shock acceleration), higher energy exponential cutoffs (log10(/V) 19.6), and an increase in the highest energy secondary photons and neutrinos. The impact of magnetic field strength on light nuclei emission is the greatest for the more modern Sibyll2.3c and EPOS-LHC hadronic interaction models.
@@
The Hackstein et al. (CLUES) ’astrophysical1R’ model [23], characterized by B = 0.064 nG and BRMS = 1.2 nG, emerged as the best fit magnetic field model for the constant nuclei fraction fits. This is compatible with a previous estimate in [51] of extragalactic magnetic fields using detected cosmic ray multiplets. Using the given Z*B*D = 30 nG*Mpc (found with a mean energy cut of E 27.7 EeV), the magnetic field of B = 0.064 nG, and the CLUES-SIBYLL best fit composition (Z 2.7) results in an average particle trajectory length of about 1.7102 Mpc. Using the energy cut E 27.7 EeV from [51] and the CLUES-SIBYLL best-fit composition, the mean trajectory length of simulated particles is 2.3102 Mpc.
@@
In conclusion, our comprehensive study provides strong evidence supporting the significant contribution of FR0 radio galaxies to the UHECR energy density. This work not only enhances our understanding of potential UHECR sources but also opens new avenues for multi-messenger astrophysics investigations.
@@ Acknowledgments
Financial support was received from the Austrian Science Fund (FWF) under grant agreement number I 4144-N27 and the Slovenian Research Agency - ARIS (project no. N1-0111). LM acknowledges support from the DFG within the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters - From Source to Signal”. MB has for this project received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 847476. GB acknowledges financial support from the State Agency for Research of the Spanish MCIU through the ”Center of Excellence Severo Ochoa” award to the Instituto de Astrofísica de Andalucía (SEV-2017-0709) and from the Spanish ”Ministerio de Ciencia e Innovación” (MICINN) through grant PID2019-107847RB-C44. The views and opinions expressed herein do not necessarily reflect those of the European Commission. The Slovenian National Supercomputing Network provided extensive computing support.
Appendix A
In this appendix are figures for intermediate results of the evolving fraction fits. Figure 14 shows the resulting after propagation for each emitted nuclei species for all ten configurations of magnetic field and EAS model. These are the values used to fit the composition data and have some dependence on the energy spectrum parameters. Figure 15 shows the 44 free composition observed fraction parameters that were fit to the data. The final emission fractions (Figure 11) are the average of the observed fractions in emitted energy bins with the ratio of observed particles to those emitted taken into account (Shown in Figure 16).

### v2

@@ 6.2 Evolving-Fraction Fits
The evolving-fraction combined-fit results for all five simulated intergalactic media and two EAS models
@@
(EPOS-LHC and Sibyll2.3c) compared to Auger data [27, 28] are shown in Figure 10. The QGSJetII-04 model was excluded from the evolving-fraction analysis. The mean log mass number shown in Figure 10(a) illustrates that the inclusion of 44 nuclei fractions leads to a nearly perfect fit across all configurations due to the large number of parameters.
@@
However, similar to the constant-fraction fits 1 nG random fields results having a better fit to the lower energy bins, these fields are the best fit to the two highest-energy heavier bins with the evolving fractions. As with the constant-fraction fits only the Rand.B configurations sufficiently account for the highest-energy bin of the energy spectra as shown in Figure 10(b). Continuing the trend observed in constant-fraction fits, the superior performance of the longer correlation length 1 nG (Rand.B) field at the highest energies remains surprising, challenging the assumption that FR0s are minor contributors at extreme energies. And again, the CLUES model essentially does not contribute to the highest energy bin.
@@
The best-fit parameters are presented in Table 5 and Figure 11, along with corresponding 1 equivalent confidence intervals around the best fit. Table 6 shows the most-probable value of the bootstrap sample distributions using an four-dimensional multivariate kernel density estimation (KDE) and the 1 standard deviations of the KDEs. A 48-dimensional KDE estimation has not been done due to time and computing constraints and the 4D KDE should be a reasonable approximation.
@@
For each EAS model and magnetic field, Table 5 and Table 6 show the total , emission spectral index (), exponential rigidity-dependent cutoff (log10(R/V)), particle trajectory cutoff (D), and relative energy-spectrum normalization (). Figure 11 shows the fitted FR0 emission elemental fractions as a function of energy for all 10 configurations.
@@
As indicated in Table 5, similar to findings from constant-fraction fits, the CLUES, Rand.A, and Rand.B magnetic fields consistently deliver the best fit results. The Rand.B field, in particular, does stand out as the optimal fit for both the EPOS-LHC and Sibyll2.3c models, with the Rand.-B-SIBYLL configuration achieving the overall best fit (Rand.B-EPOS is a very close second). EPOS-LHC is the best fit for all magnetic field models except Rand.B-EPOS but the difference from SIBYLL is very small. The second-best magnetic field after Rand.B is CLUES.
6.2.1 Energy-Spectrum Parameters
@@
Overall, the evolving-fraction fit energy-spectrum parameters of Table 5 and Table 6 again indicate that both magnetic field strength B, and RMS BRMS, have an impact on the expected cosmic-ray source emission energy spectrum. Generally, the emission spectral index, E-γ, increases with the magnetic field strength B. However, the exception is the Rand.B field model where this trend does not apply. The exponential rigidity-dependent cutoff, R, generally rises with an increase in BRMS, except in the case of the Rand.A field model. Furthermore, for smaller magnetic fields, R decreases with a heavier composition, while for larger fields, this trend is reversed, demonstrating the complex interaction between magnetic field strength and particle composition. The trajectory cutoff, D, is fairly stable across most fields, with the notable exception of the Dolag field which has a significantly higher BRMS of 11 nG.
@@
Despite the no-field and Dolag configurations consistently showing poorer fits, similar to the constant-fraction scenario, the inclusion of a large number of parameters in the evolving-fraction fits leads to excellent overall fit quality for all configurations. Therefore, the bounds found for the best fit of these parameters from Table 2, using all configurations, are:
@@
Note, that these bounds are larger than the constant-fraction fit bounds due the inclusion of the no-field and Dolag configurations.
@@
For the bootstrap KDE fit most-probable parameters from Table 6 the ranges found are:
@@
Compared to the recent Auger result of [37] these parameter values are similar to the reported second minima of their models and the low- source distance model ( 0.02).
@@ 6.2.2 UHECR Composition Parameters
The FR0 emission fractions for the evolving energy-dependent (observed) fraction fit of proton, helium, nitrogen, silicon, and iron emitted nuclei are shown in Figure 11. Similar to the constant-fraction fits, the emission fractions of protons and helium exhibit instability, with dominance of each varying due to propagation effects, as shown in Figure 5. The combined fraction of light components () demonstrates greater stability across different configurations as shown in Figure 11(f). Interestingly, for all configurations the fraction of emitted protons increases with energy above 1019 eV as shown in Figure 11(a).
@@
The ranges found for the energy-bin average FR0-emission best-fit elemental fractions (in %) from Figure 11 are:
@@
6.2.3 Multimessenger Photons and Neutrinos
The evolving-fraction fit cosmogenic integral photon and all-flavor neutrino spectra are shown in Figures 12(a) and 12(b). These results are for the best-fit configuration using the Rand.B 1 nG random magnetic field and Sibyll2.3c-based mass-composition, compared to the no-field scenario. Although the Sibyll2.3c hadronic interaction model reconstructs a heavier mass for data relative to other models, the evolving-fraction fit still predicts that approximately 66% of the FR0 emission consists of light nuclei.
At lower energies (up to eV), the integral photon flux is compatible with pure proton cosmogenic predictions, shifting towards a theoretical mixed composition at higher energies. Interestingly, the neutrino flux resembles predictions for a pure iron prediction, which may be partially influenced by the simulation constraint of relatively close FR0 sources ( 0.2). Given the wide range of theoretical predictions and current experimental upper limits the FR0 simulated flux is reasonable. Overall, for both photon and neutrino spectra, it can be seen that a magnetic field results in a larger flux, particularly at higher energies, as shown in Figure 13.
@@ 7 Summary & Conclusions
This work presents an in-depth study of FR0 radio galaxies, a prevalent but less luminous class of jetted active galaxies, and their potential emission features as ultra-high-energy cosmic-ray (UHECR) sources. Although individually less luminous than more powerful radio galaxies, the abundance of FR0 galaxies may render them significant contributors to the UHECR energy density observed on Earth.
@@
To investigate this hypothesis, extensive CRPropa3 [26] simulations were conducted, modeling isotropic cosmic-ray emissions extrapolated from measured FR0 properties (density and luminosity distributions and luminosity-redshift source evolution). These simulations were then fit to the mass-composition (expressed in mean log mass number ) and energy spectra of UHECR data from the Pierre Auger Observatory [27, 28].
@@
The results from these combined fits, utilizing two fit models, three EAS models, and five intergalactic magnetic field configurations (both random and structured), indicate that emissions from FR0 galaxies can effectively account for the observed UHECR energy spectrum and composition. The overall compatibility, in addition to the results of our previous study of the capability of FR0 radio galaxies to accelerate cosmic rays up to ultra-high energies [6], supports the viability of FR0s as significant contributors to the observed UHECR flux. Furthermore, the secondary photon and neutrino fluxes predicted from UHECR interactions are well within current observational limits, emphasizing the potential of future multimessenger observations to further refine our understanding of FR0 galaxies as UHECR sources.
@@
For the most extreme energy bin, the configurations tested are generally below the expected flux (only Dolag and longer correlation length 1 nG Rand.B magnetic field models have overlapping error bars, as shown in Figure 6(b)). Additionally, over all configurations, there is a small underfitting of the last energy bin of (only Dolag and the smaller correlation length 1 nG Rand.A magnetic field models are outside the error bars as shown in Figure 6(a)). These two results may allow the possibility of a different extreme-energy heavy nuclei source type to become the dominant UHECR flux contributor.
@@
These findings underscore previous results that show the importance of considering the effects of intergalactic magnetic fields on UHECR propagation (such as [39]). Results across the different magnetic field models tested indicate their significant influence on the arrival energy spectrum and composition of cosmic rays from FR0 galaxies, as well as up to orders-of-magnitude variations on the secondary photon and neutrino spectra. Larger magnetic fields result in better fits and a need for a greater emission of nuclei heavier than helium - additionally, generally, slightly harder emission spectra ( 2.4 compatible with Fermi shock acceleration), higher energy exponential cutoffs (log10(R/V) 19.6), and an increase in the highest energy secondary photons and neutrinos. The impact of magnetic field strength on light-nuclei emission is the greatest for the more modern Sibyll2.3c and EPOS-LHC hadronic interaction models.
@@
The Hackstein et al. (CLUES) ’astrophysical1R’ model [23], characterized by B = 0.064 nG and BRMS = 1.2 nG, emerged as the best-fit magnetic field model for the constant nuclei fraction fits. This is compatible with a previous estimate in [52] of extragalactic magnetic fields using detected cosmic-ray multiplets. Using the given Z*B*D = 30 nG*Mpc, the magnetic field of B = 0.064 nG, and the CLUES-SIBYLL best-fit composition (Z 2.7) results in an average particle trajectory length of about 1.7102 Mpc. Using the mean energy cut E 27.7 EeV from [52] and the CLUES-SIBYLL best-fit composition, the mean trajectory length of simulated particles is
2.3102 Mpc.
@@
In conclusion, our comprehensive study provides strong evidence supporting the significant contribution of FR0 radio galaxies to the UHECR energy density. This work not only enhances our understanding of potential UHECR sources but also opens new avenues for multimessenger astrophysics investigations.
@@ Acknowledgments
Financial support was received from the Austrian Science Fund (FWF) under grant agreement number I 4144-N27 and the Slovenian Research Agency—ARIS (project no. N1-0111). LM acknowledges support from the DFG within the Collaborative Research Center SFB1491 ”Cosmic Interacting Matters—From Source to Signal”. MB has for this project received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Sklodowska-Curie grant agreement No 847476. GB acknowledges financial support from the State Agency for Research of the Spanish MCIU through the ”Center of Excellence Severo Ochoa” award to the Instituto de Astrofísica de Andalucía (SEV-2017-0709) and from the Spanish ”Ministerio de Ciencia e Innovación” (MICINN) through grant PID2019-107847RB-C44. The views and opinions expressed herein do not necessarily reflect those of the European Commission. The Slovenian National Supercomputing Network provided extensive computing support.
Appendix A Evolving Fraction Fit
In this appendix are figures for intermediate results of the evolving-fraction fits. Figure 14 shows the resulting after propagation for each emitted nuclei species for all ten configurations of magnetic field and EAS model. These are the values used to fit the composition data and have some dependence on the energy-spectrum parameters. Figure 15 shows the 44 free composition observed fraction parameters that were fit to the data. The final emission fractions (Figure 11) are the average of the observed fractions in emitted-energy bins with the ratio of observed particles to those emitted taken into account (Shown in Figure 16). Each figure in this appendix has two legends—the first emphasizes figure element shapes, while the second includes marker colors for all fits, where each base color represents a magnetic field model.